# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USAthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USA

### v1

@@ Appendix B CZ gate on-off ratio
Appendix C Single-qubit characterization
C.1 1Q gate characterization
Single-qubit gates are realized by four wave mixing parametric beamsplitter-type interactions implemented by a SQUID-transmon coupler [35]. We apply a flux drive to the SQUID coupler at half the frequency detuning between the two cavities of a dual-rail qubit to enable a parametric interaction via the Hamiltonian $\mathcal{H}/\hbar=\frac{g_{\mathrm{bs}}}{2}\left(\hat{a}^{\dagger}\hat{b}+\hat{a}\hat{b}^{\dagger}\right)$. To characterize single-qubit gates, we perform both individual and simultaneous RB. For these measurements we use a gateset consisting of $\{X_{\frac{\pi}{2}},Z,Z_{\pm\frac{\pi}{2}}\}$ from which we generate the 24 single-qubit Cliffords. For each Clifford sequence length we average over 20 different randomized sequences. Performing RB individually on control and target qubit, we extract a gate fidelity in excess of $>99.98\%$ with a small decrease in postselected fraction of $<0.1\%$ per Clifford detected by our end-of-the-line measurement (see Fig. 7). When operated simultaneously, the control and target qubit fidelities are observed to modestly decrease by $0.001\%$ and $0.002\%$, respectively for the control and target qubits. We attribute this effect to a small measurable cross-Kerr between two dual-rail qubits ($a_{2}$ and $b_{1}$ cavities) of $\mathrm{6.64\>kHz}$. Compared to our single-qubit beamsplitter rates of $1.78\mathrm{\>MHz\>}(2.55\mathrm{\>MHz})$ for the control (target) qubit we estimate simultaneous operation infidelities to be limited by $(\chi_{ab}/g_{aa(bb)})^{2}\sim 1.3\times 10^{-5}$.
C.2 SPAM
Dual-rail measurement are implemented by performing simultaneous single-shot cavity measurements that determine whether a photon is present in each cavity individually, extracting whether the system is in |$\ket{01}$ or |$\ket{10}$. This measurement is designed to also have an important property: it can detect for both photon loss and also first-order errors, like decoherence or transmon readout misassignments, during the measurement itself. In addition, this measurement can be repeated for multiple rounds and correlating the results can further suppress misassignment errors. This postselection capability comes at the cost of a first-order decrease in postselected fraction, however, we benefit from a exponential decrease in missassignments, which is highly beneficial for characterization of our CZ gate. Additional details can be found in ref [33]. We benchmark performance for state preparation and measurement (SPAM) and extract logical misassignment errors of $0.023\%$ and $0.020\%$ for the target and control dual-rail qubit respectively. The overall SPAM erasure errors, defined to be assigned erasure outcomes when attempting to prepare in either |$\ket{01}$ or |$\ket{10}$, are measured to be $6.79\%$ and $6.43\%$ for the target and control qubit. Finally we benchmark the leakage detection properties by initializing the dual-rail qubit in the |$\ket{00}$ and we observe erroneous codespace assignments to be $0.39\%$ and $0.34\%$ for the target and control qubit.
Appendix D Computing fidelities
@@
with $d=2^{N}=4$, $p_{i,r}$ are the depolarizing parameter for interleaved and reference RB, respectively.
@@
Appendix E Leakage propagation
@@
Table 4: Leakage propogation properties for CZ gate
E.1 Leakage before a CZ gate
@@
E.2 Leakage during a CZ gate
@@
$\Lambda_{\mathrm{leakage}}(\rho)=\frac{1}{2}\left(\rho+\mathrm{CZ}\rho\mathrm{CZ}\right)+\frac{4i}{3\pi}\left(\mathrm{CZ}\rho-\rho\mathrm{CZ}\right),$ (12)
@@
In general, only the diagonal terms will survive a Pauli twirling procedure. In our case, Pauli twirling results in the simple effective error channel
$\Lambda_{\mathrm{leakage}}(\rho)=\frac{1}{2}\left(\rho+\mathrm{CZ}\rho\mathrm{CZ}\right).$ (13)
In an error correction context, this twirling occurs naturally from the Hadamard gates interleaved between the CZ gates. In our experiments, this effective twirling arises from our RB sequences, or from the single $X$ gate in our quantum state tomography sequence. One may notice that a channel of the form $\mathrm{CZ}\rho\mathrm{CZ}$ is not a Pauli channel, but a Clifford channel. Pauli twirling does not remove the highly correlated errors between the control and target qubit. Preserving this correlation is essential for enabling delayed erasure detection when using this gate to perform stabilizer measurements.
@@
E.3 Analysis of experimental results

### v2

@@ Appendix A Single-qubit characterization
Single-qubit gates are realized by four wave mixing parametric beamsplitter-type interactions implemented by a SQUID-transmon coupler [36]. We apply a flux drive to the SQUID coupler at half the frequency detuning between the two cavities ($a$ and $b$) of a dual-rail qubit to enable a parametric interaction via the Hamiltonian $\mathcal{H}/\hbar=\frac{g_{\mathrm{bs}}}{2}\left(\hat{a}^{\dagger}\hat{b}+\hat{a}\hat{b}^{\dagger}\right)$. To characterize single-qubit gates, we perform both individual and simultaneous RB. For these measurements we use a gateset consisting of $\{X_{\frac{\pi}{2}},Z,Z_{\pm\frac{\pi}{2}}\}$ from which we generate the 24 single-qubit Cliffords. For each Clifford sequence length we average over 20 different randomized sequences. Performing RB individually on control and target qubit, we extract a gate fidelity in excess of $>99.98\%$ with a small decrease in postselected fraction of $<0.1\%$ per Clifford detected by our end-of-the-line measurement (see Fig. 5). When operated simultaneously, the control and target qubit fidelities are observed to modestly decrease by $0.001\%$ and $0.002\%$, respectively for the control and target qubits. We attribute this effect to a small measurable cross-Kerr between two dual-rail qubits ($a_{2}$ and $b_{1}$ cavities) of $\mathrm{6.64\>kHz}$. Compared to our single-qubit beamsplitter rates of $1.78\mathrm{\>MHz\>}(2.55\mathrm{\>MHz})$ for the control (target) qubit we estimate simultaneous operation infidelities to be limited by $(\chi_{ab}/g_{aa(bb)})^{2}\sim 1.3\times 10^{-5}$.
A.2 SPAM
Dual-rail measurement are implemented by performing simultaneous single-shot cavity measurements that determine whether a photon is present in each cavity individually, extracting whether the system is in |$\ket{01}$ or |$\ket{10}$. This measurement is designed to also have an important property: it can detect for both cavity photon loss and also can flag hardware errors during the measurement protocol, like decoherence or transmon readout misassignments, during the measurement itself. Detailed characterization of dual-rail SPAM are shown in Table 3. In addition, this measurement can be repeated for multiple rounds and correlating the results can further suppress misassignment errors. This postselection capability comes at the cost of a first-order decrease in postselected fraction, however, we benefit from a exponential decrease in missassignments, which is highly beneficial for characterization of our CZ gate. Additional details can be found in ref [34]. We benchmark performance for state preparation and measurement (SPAM) and extract logical misassignment errors of $0.023\%$ and $0.020\%$ for the target and control dual-rail qubit respectively. The overall SPAM erasure errors, defined to be assigned erasure outcomes when attempting to prepare in either |$\ket{01}$ or |$\ket{10}$, are measured to be $6.79\%$ and $6.43\%$ for the target and control qubit. Finally we benchmark the leakage detection properties by initializing the dual-rail qubit in the |$\ket{00}$ and we observe erroneous codespace assignments to be $0.39\%$ and $0.34\%$ for the target and control qubit.
Appendix B Two-qubit gate calibration
In this section, we discuss the calibration of the CZ gate in detail. The calibration procedure is illustrated in the flowchart in Fig. 6a, which starts with calibrating the photon swap operation between the control qubit cavity $a_{2}$ and the coupler $c$, referred to as the cavity-coupler swap. The underlying interaction for this cavity-coupler swap is identical to that used for the single-qubit gate. However, in this case, the parametric flux pump is applied at half the frequency difference of the Stark shifted coupler $c$ and cavity $a_{2}$. A coherent transfer of a single photon between the cavity and the coupler is shown in Fig. 6b where we also present the stroboscopic time evolution of the photon population in the cavity $P_{\rm{cav}}$ at the resonant pump frequency, which follows
$P_{\rm{cav}}(t)=\frac{1}{2}e^{-{\kappa t}}\left(1+e^{-\kappa_{\phi}t}{\rm{cos}}(g_{ac}t)\right).$ (2)
Here, $g_{ac}$ represents the rate of the cavity-coupler swap, while $\kappa$ and $\kappa_{\phi}$ denote the photon decay rate and the dephasing rate when the pump is turned on. As the first step in calibrating the cavity coupler swap operation, we find the amplitude $A_{p}$ that maximizes the metric $g_{ac}/\kappa_{\phi}$, which leads to the highest fidelity of the cavity-coupler swap operation when you postselect for photon loss. Following this, we calibrate the duration of the swap pulse $t_{\mathrm{SWAP}}$ by repeatedly applying the pulse after starting with a photon in cavity $a_{2}$ and minimizing the population in the cavity $a_{2}$ at the end of odd number of swap operations (see Fig. 6c). When the swap operation is used in the SWS gate sequence, an additional parameter needs to be specified which is the relative phase difference between the first and the second swap operations. To understand why this is important, we first note that at the end of the CZ gate we want all the population in cavity $a_{2}$ to return to its original value irrespective of which two qubit state we start in. This condition is trivially satisfied for starting states |$\ket{10g01},\ket{10g10}$ and |$\ket{01g01}$ since in these cases either no transfer of population takes place between cavity $a_{2}$ and coupler $c$ or an on-resonant swap takes place, either way the phase of the second swap operation is irrelevant. The only non-trivial case occurs when we start in the state |$\ket{01g10}$, in which case we perform a swap operation detuned by $\chi_{bc}$ (see Fig. 12). In this case we can close the Bloch sphere trajectory by choosing the correct phase of the second swap operation in the gate. To calibrate this parameter we initialize our system in the state |$\ket{01g10}$ and apply the SWS gate sequence while sweeping the phase of the second swap pulse and measure the total erasure fraction of the circuit. We choose the swap back phase $\phi_{\mathrm{swap}}$ that minimizes the erasure fraction (See Fig. 6d). Since our gate is excitation-preserving, every time we leave population in the coupler mode $c$ we will find the control qubit in the vacuum state. Therefore, minimizing the erasure fraction is equivalent to maximizing the population in cavity $a_{2}$ at the end of the gate. It is important to note that as the delay time varies, the swap back phase should be calibrated and adjusted accordingly prior to measuring the corresponding entangling phase $\phi_{e}$ as illustrated in the flowchart, until the delay time is accurately calibrated.
Figure 6: Calibration of CZ gate. a Flowchart of the CZ gate calibration procedure. b Left: Schematic illustration of swap between the control qubit cavity $a_{2}$ (red) and coupler $c$ (purple). Right: Photon population in the control qubit cavity as a function of parametric pump frequency $\omega_{p}$ and duration of the pump, and additionally the stroboscopic evolution at the resonant pump frequency marked by a red dashed line. c-f. Each Fig. shows the pulse sequence and the corresponding calibration result. c. Calibration of duration of cavity-coupler swap by measuring the population in the control qubit cavity after odd numbers of swap. d. Calibration of the phase of rotation for the second swap operation in the CZ gate by minimizing the erasure fraction. e. Calibration of the delay time of the CZ gate by measuring the entangling phase per CZ gate, which should corresponds to $\pi$ for the calibrated delay time marked by dashed line. f. Calibration of single-qubit Z rotations per CZ gate. The Ramsey-like protocol should be performed on either the control or target qubit to measure Z rotation of each qubit.
With the cavity-coupler swap time and the swap back phase for a given delay time $t_{\mathrm{wait}}$ calibrated, the unitary of the SWS gate can be written as the following,
where $\phi_{e}$ is the entangling phase, $\phi_{a}$ and $\phi_{b}$ are respective $Z$-phase on each qubit accumulated during the gate. Next, to realize CZ unitary, we calibrate the wait time $t_{\mathrm{wait}}$ such that $\phi_{e}=\pi$. For cavity-coupler swaps with instantaneous pulse rise times and zero cavity-cavity cross-Kerr $\chi_{ab}$ between the cavities $a_{2}$ and $b_{1}$, the wait time can be analytically calculated (see Eq. 56). However, in the case of non-zero pulse rise time and $\chi_{ab}$, we need to experimentally determine the wait time. To measure $\phi_{e}$ for a given wait time, we perform a Ramsey-like sequence [47] on either the control or the target qubit . In this protocol, we start with both qubits in the state |$\ket{0_{L}}$ and apply a $\sqrt{X}$ gate to one of the qubits to prepare a superposition state $\frac{1}{\sqrt{2}}(|0_{L}0_{L}\rangle-i|1_{L}0_{L}\rangle)$. Following that we repeatedly apply $N$ CZ gates interrupted by $X$ gates on both qubits in the middle. This protocol then creates the state $\frac{1}{\sqrt{2}}(e^{iN\phi_{e}/2}|1_{L}1_{L}\rangle-i|0_{L}1_{L}\rangle)$ where the phase $N\phi_{e}/2$ is accumulated as a relative phase shift between the states and the phases $\phi_{a}$ and $\phi_{b}$ are factored out as global phases. The entangling phase $\phi_{e}$ measured as a function of wait time $t_{\mathrm{wait}}$ and number of repeats is shown in Fig. 6e from which we find the optimal wait time that leads to $\phi_{e}=\pi$.
Once the delay time is precisely calibrated, the last two parameters to be calibrated are the $Z$-phase rotation for the control and target qubit $\phi_{a}$ and $\phi_{b}$. These can also be measured via Ramsey-like protocol on the control and target qubit in similar manner as the entangling phase $\phi_{e}$. The only difference is that we do not perform $X$ gate in the middle of the sequence as the echo pulses will convert the phases $\phi_{a}$ and $\phi_{b}$ that we are interested in measuring into in to global phases. Fig. 6f shows these local Z-phases as a function of number of repeated CZ gates for the wait time corresponding to entangling phase $\phi_{e}=\pi$ where the extracted slope of each trace corresponds to the phases $\phi_{a}$ and $\phi_{b}$ accumulated per gate.
Appendix C CZ gate on-off ratio
@@
Appendix D Inferring gate fidelities
@@
with $d=2^{N_{q}}=4$, $p_{i,r}$ are the depolarizing parameter for interleaved and reference RB, respectively.
@@
Appendix E Evaluating the accuracy of IRB with postselection
Figure 8: $\mathrm{CZ}$ gate infidelity inferred from simulated IRB experiments vs. the infidelity computed from Eq. 10. 500 random dephasing and leakage rates are chosen uniformly distributed with $p^{\text{control}}_{Z},\,p^{\text{target}}_{Z}\in[1\times 10^{-4},2\times 10^{-3}]$, $p_{ZZ}\in[5\times 10^{-5},2\times 10^{-4}]$, and $p^{\text{control}}_{\text{leakage}},\,p^{\text{target}}_{\text{leakage}}\in[1\times 10^{-4},4\times 10^{-3}]$. Each IRB fidelity simulation uses the exact same random gate sequence as the measurement in Fig. 2. The solid lines guide the eye to convert our measured $\mathrm{CZ}$ infidelity to the infidelity when error detection immediately follows the gate.
Appendix D utilizes IRB to estimate $\mathrm{CZ}$ gate fidelities at short depths with a linear approximation of the decay exponential. Here, we demonstrate that the experimental results are a reliable estimator for the true underlying gate errors. The addition of error detection to RB is novel and, to our knowledge, only utilized here and in neutral atom platforms [27]. These error detection measurements are high-fidelity [34], but leakage misassignment errors do occur at $10^{-4}$ rates. Additionally, the infidelity of our $\mathrm{CZ}$ gate exceeds those of single qubit rotations by more than one order of magnitude. This produces an asymmetry between some Cliffords and the interleaved $\mathrm{CZ}$ gate that can negatively impact the $\mathrm{CZ}$ fidelity estimate [49]. Between the addition of error detection, gate asymmetry, and limited circuit depth, it is valid to worry that IRB may not provide an accurate estimate of $\mathrm{CZ}$ gate fidelity. In this section, we define a parametrized channel for our $\mathrm{CZ}$ gate, add postselection to the standard definition of average gate fidelity [50], and simulate IRB under different $\mathrm{CZ}$ error rates to evaluate its estimates of the postselected gate fidelity.
First, we discuss a simplified leakage model for a DR qubit. For a single DR, we represent the logical and leakage states with qutrit states |$\ket{0}=\ket{0_{L}}$, |$\ket{1}=\ket{1_{L}}$, |$\ket{2}=\ket{*_{L}}$. Thus, for each $x\in\{\mathrm{target},\;\mathrm{control}\}$ in a two DR system, the leakage channel is:
$\Lambda^{x}_{\text{leakage}}(\rho)$ $=(1-p^{x}_{\text{leakage}})\rho+p^{x}_{\text{leakage}}K^{x}\rho(K^{x})^{\dagger}$ (7)
[…]
Detectable leakage errors form the dominant component of the $\mathrm{CZ}$ error channel, but residual dephasing arises from intrinsic sources and control qubit leakage during gates. To form the dephasing error channel, note that dephasing in the computational subspace {|$\{\ket{0_{L}},\ket{1_{L}}\}$ is generated by the $\sigma_{z}$-like Gell-Mann matrix $\lambda_{3}$ [51] with non-zero diagonal $(1,-1,0)$. The resulting dephasing Kraus operator on a single qutrit is $K_{3}=\exp{(-i\pi\lambda_{3}/2)}$ ). This produces the two qutrit dephasing channel:
$\Lambda_{\text{dephas.}}(\rho)$ $=(1-p^{\text{control}}_{Z}-p^{\text{target}}_{Z}-p_{ZZ})\rho$ (8)
$\;\;+p^{\text{control}}_{Z}K_{3}^{\text{control}}\rho(K_{3}^{\text{control}})^{\dagger}$
$\;\;+p^{\text{target}}_{Z}K_{3}^{\text{target}}\rho(K_{3}^{\text{target}})^{\dagger}$
$\;\;+p_{ZZ}(K_{3}^{\text{control}}\otimes K_{3}^{\text{target}})\rho(K_{3}^{\text{control}}\otimes K_{3}^{\text{target}})^{\dagger}.$
The $Z$ subscripts serve as a reminder that dephasing occurs in the qubit subspace.
Finally, we can write a simple $\mathrm{CZ}$ channel $\mathcal{E}$ as the composition of leakage and dephasing channels
$\mathcal{E}(\rho)$ $=\Lambda^{\text{control}}_{\text{leakage}}\circ\Lambda^{\text{target}}_{\text{leakage}}\circ\Lambda_{\text{dephase}}(\mathrm{CZ}\rho\mathrm{CZ}^{\dagger}).$ (9)
This channel captures target qubit dephasing resulting from control qubit leakage during the $\mathrm{CZ}$ gate, albeit more simply than the detailed picture presented in Appendix F. The rates parametrizing each individual error channel are derived from the error-per-gate results in Fig. 3.
Our RB circuits only include leakage detection at the end of a circuit, unlike other approaches that interleave checks with batches of gates [27, 32]. Here we explain why this technique is valid for our system. DR qubit seepage via cavity photon heating occurs at a rate at least 1000 times slower than leakage [34], obviating the need for mid-circuit checks here or other more sophisticated techniques [52, 53]. This means that the fidelity of a standalone gate immediately followed by leakage detection should be nearly equivalent to the fidelity of a gate with delayed leakage detection. An RB experiment with postselection should then infer the fidelity of a postselected gate, up to misassignment errors in the leakage detection. This motivates a modification to the standard entanglement fidelity definition found in [50]:
$F_{e}(\mathcal{E},\mathrm{CZ})=\sum_{j,k}\frac{\alpha_{jk}\mathrm{Tr}(\mathrm{CZ}U_{j}\mathrm{CZ}(M\mathcal{E}(\rho_{k})M^{\dagger}))}{d^{3}\mathrm{Tr}(M\mathcal{E}(\rho_{k})M^{\dagger})}.$ (10)
The $U_{j}=\sum_{k}\alpha_{jk}\rho_{k}$ form a unitary basis on the two qubit space [50, 54] with pure state representation $\rho_{k}$. Using a state basis here is essential for postselection after propagation through the $\mathrm{CZ}$ channel $\mathcal{E}$ as the $U_{j}$ can be traceless. The postselection measurement operator $M$ projects into the two-qubit codespace, with misassignments characterized in Table 3 that appropriately penalize the gate fidelity [34]. The denominator in Eq. 10 renormalizes the fidelity contribution of each $\rho_{k}$ by its postselected fraction. This is similar to the approach taken in [55], where the entanglement fidelity is divided by the average postselected fraction with the intent to preserve channel linearity.
Now that we can compute the postselected fidelity of our $\mathrm{CZ}$ channel, we compare this to the fidelity inferred from IRB simulations across a range of $\mathrm{CZ}$ error rates. We follow a standard technique used in [50, 48, 49] using the superoperator formalism to propagate a density matrix through a random sequence of gates. There are two more gates required to generate the Clifford group: $X_{\frac{\pi}{2}}$ and virtual-$Z$. We compute the superoperator implementing both individual and simultaneous $X_{\frac{\pi}{2}}$ operations using the coherences and gate durations in Tables 2 and 3. It is important to include the residual $a_{1}-b_{2}$ cross-Kerr to capture unitary gate errors. Virtual-$Z$ gates are perfect unitaries in the simulation.
Having described our model for all operations, we proceed to simulate IRB and extract two-qubit gate fidelities. All of our simulations implement the same random gate sequence used for the measurement in Fig. 2. Due to finite sampling of random sequences (20 at each depth), we expect noise in the inferred infidelity particular to each sequence. Using error rates centered around estimated values from Fig. 3, we find that IRB systematically underestimates infidelities across the range, including by $25\%$ at our measured value (marked on Fig. 8). In absolute terms, this only adjusts the infidelity from $0.1\%$ up to $0.13\%$, confirming a high-fidelity gate. Other investigations have found comparable fractional errors at similarly high gate fidelities, albeit under maximally deleterious unitary errors [49]. The underestimate persists even when using perfect leakage detection measurements in the simulation, suggesting that the bias may be a product of this particular set of random gates. Simulations of other random sequences produce different slopes and offsets such as $0.984(1)$ and $-0.000150(2)$, respectively, supporting this claim.
Appendix F Leakage propagation
@@
Table 4: Leakage propagation properties for CZ gate
F.1 Leakage before a CZ gate
@@
F.2 Leakage during a CZ gate
@@
$\Lambda_{\mathrm{leakage}}(\rho)=\frac{1}{2}\left(\mathds{1}\rho\mathds{1}+\mathrm{CZ}\rho\mathrm{CZ}\right)+\frac{4i}{3\pi}\left(\mathrm{CZ}\rho\mathds{1}-\mathds{1}\rho\mathrm{CZ}\right),$ (15)
@@
In an error correction context, measuring stabilizers will “digitize” our quantum errors and the gate’s leakage error channel can be effectively modeled by the following effective error channel where the imaginary terms have been removed:
$\Lambda_{\mathrm{CZ}}(\rho)=\frac{1}{2}\left(\mathds{1}\rho\mathds{1}+\mathrm{CZ}\rho\mathrm{CZ}\right).$ (16)
Terms in the channel of the form ‘$\mathrm{CZ}\rho\mathds{1}$’ and ‘$\mathds{1}\rho\mathrm{CZ}$’ are annihilated when we perform the projective stabilizer measurements during quantum error correction with any stabilizer code. One may take notice that Eq. 16 describes a Clifford error channel, not a Pauli error channel. We must be careful not to apply any active Pauli twirling procedures to our qubits, since this would diagonalize our error channel and reduce the error correlations between our qubits. Preserving this correlation is essential for enabling delayed erasure detection when using this gate to perform stabilizer measurements in a surface code, since it ensures our leakage errors do not lead to ‘bad’ hook errors which can reduce the effective distance of the code.
@@
[…]
$\mathcal{E}(\rho)=\Lambda^{\text{control}}_{\text{leakage}}\circ\Lambda^{\text{target}}_{\text{leakage}}\circ\Lambda_{\text{dephase}}(\mathrm{CZ}\rho\mathrm{CZ}^{\dagger}),$ (17)
with the following channel definitions:
$\Lambda^{\text{control}}_{\text{leakage}}(\rho)$ $=(1-p_{e}^{c})\mathds{1}\rho\mathds{1}+p_{e}^{c}\hat{L}_{c}\Lambda_{\mathrm{CZ}}(\rho)\hat{L}_{c}^{\dagger},$ (18)
$\Lambda^{\text{target}}_{\text{leakage}}(\rho)$ $=(1-p_{e}^{t})\mathds{1}\rho\mathds{1}+p_{e}^{t}\hat{L}_{t}\Lambda_{\mathrm{CZ}}(\rho)\hat{L}_{t}^{\dagger},$ (19)
$\Lambda_{\mathrm{dephase}}(\rho)$ $=(1-p_{Z}^{c}-p_{Z}^{t}-p_{ZZ})\mathds{1}\rho\mathds{1}$ (20)
$+p_{Z}^{c}Z_{c}\rho Z_{c}+p_{Z}^{t}Z_{t}\rho Z_{t}$ (21)
$+p_{ZZ}Z_{t}Z_{c}\rho Z_{c}Z_{t}.$ (22)
For this channel definition we define $p_{e}^{c(t)}$ to be the probability of erasure on the control (target) qubit during a CZ gate, and $p_{Z}^{c(t)}$ to be the probability of a phase flip error on the control (target) qubit. Lastly, $p_{ZZ}$ is the probability of a correlated dephasing error that affects both qubits.
F.3 Analysis of experimental results